# Generating Directed Graphs with Dual Attention and Asymmetric Encoding

### v2

@@ Appendix C Sampling Optimization in Directo
Together, these sampling parameters constitute a flexible, principled toolkit for tailoring the sampling procedure to diverse graph domains and structural requirements. In our case, we optimize these parameters at sampling time to improve the performance of Directo. In particular, we individually optimize each of them for the main results, and perform the final sampling using the best performing parameters in terms of V.U.N (a detailed description is available in Appendix G.2). and ratio performance. A study of the effect of these parameters on sampling performance can be found in H.8.
@@ F.1 Validity metrics
Given a graph $G=(V,E)$ with $n=|V|$ nodes and $m=|E|$ edges, we want to determine whether the graph likely originates from an Erdős–Rényi model $G(n,p)$ with edge probability $p$ using a Wald test (Held & Sabanés Bové, 2013). For that, we follow the following steps distinguishing the cases in which the graphs are DAGs or not:
@@ G.2 Training setup
To perform sampling optimization, we conducted a targeted search over the three key hyperparameters: the time distortion function, the target guidance factor ($\omega$), and the stochasticity coefficient ($\eta$). Each hyperparameter was varied independently, while the others were held at their default values (identity distortion and $\omega=\eta=0$). For computational efficiency, all searches were performed using 100 sampling steps, and each configuration was evaluated using five independent sampling runs to ensure robustness. An exception was made for the TPU Tiles dataset, where only a single sampling run was performed per configuration due to the large size of the dataset and some individual graphs.
@@
Encoding Distortion $\bm{\omega}$ $\bm{\eta}$
ER-DAG
Baseline Polydec 0.3 10
RRWP Polydec 0.1 25
MagLap Polydec 0 100
SBM
Baseline Revcos 0 0
RRWP Polyinc 0.1 50
MagLap Polydec 0.01 10
TPU Tiles
Baseline Polydec 0 0
RRWP Polyinc 0.3 10
MagLap Polydec 0.5 100
Visual Genome
Baseline Cos 0.1 0
RRWP Polydec 0.3 10
MagLap Polydec 0.1 50
@@ G.3 Resources and runtime
Table 8: Training and sampling runtimes for each dataset and positional encoding variant.
Encoding Training graphs Training time (h) Graphs sampled Sampling time (min/sample)
@@
Baseline 128 13.6 40 0.125
RRWP 128 13.5 40 0.3
MagLap 128 14.4 40 0.8
@@
Baseline 128 23.7 40 0.3
RRWP 128 13.7 40 0.5
MagLap 128 21 40 2.1
@@
Baseline 5040 10 40 1.1
RRWP 5040 21 40 1.1
MagLap 5040 49 40 6.7
@@
Baseline 203 7.5 63 0.9
RRWP 203 9 63 1.1
MagLap 203 10 63 1.3
@@ H.2 Extended results for real-world datasets
Table 11: Directed graph generation performance on node-label distributional alignment metrics for the two real-world datasets. Results are presented as mean $\pm$ standard deviation across five sampling runs.
@@ H.3 The role of dual attention
Table 12 presents ablation results highlighting the impact of the dual attention mechanism in our method. The results are presented for two datasets: ER-DAG and SBM.
@@
RRWP 0.0036 $\pm$ 0.0009 0.0034 $\pm$ 0.0008 0.1276 $\pm$ 0.0065 0.0345 $\pm$ 0.0028 0.0961 $\pm$ 0.0071 21.4 $\pm$ 1.6 13.5 $\pm$ 5.4 100 $\pm$ 0.0 100 $\pm$ 0.0 13.5 $\pm$ 5.4
RRWP - Dual 0.0038 $\pm$ 0.0011 0.0037 $\pm$ 0.0012 0.0492 $\pm$ 0.0041 0.0035 $\pm$ 0.0007 0.0015 $\pm$ 0.0004 1.4 $\pm$ 0.2 87.0 $\pm$ 5.1 100 $\pm$ 0.0 100 $\pm$ 0.0 87.0 $\pm$ 5.1
MagLap 0.0033 $\pm$ 0.0008 0.0034 $\pm$ 0.0008 0.1356 $\pm$ 0.0118 0.0481 $\pm$ 0.0022 0.1250 $\pm$ 0.0017 27.7 $\pm$ 4.3 21.5 $\pm$ 2.0 100 $\pm$ 0.0 100 $\pm$ 0.0 21.5 $\pm$ 2.0
@@
In particular, we observe that, across both datasets, the inclusion of the dual attention mechanism consistently improves the validity of the generated graphs and reduces structural discrepancies as measured by clustering, spectral, and wavelet distances. In the ER-DAG setting, models with dual attention (e.g., RRWP-Dual and MagLap-Dual) significantly outperform their non-dual counterparts in validity, achieving up to 94% validity with RRWP-Dual and 91% with MagLap-Dual, compared to just 67.5% and 74%, respectively. These models also demonstrate better structural fidelity, particularly in clustering and spectral metrics. A similar trend is observed in the SBM dataset, where dual attention boosts validity (e.g., RRWP-Dual: 87%, MagLap-Dual: 77%). These results demonstrate how incorporating bidirectional information flow into the attention mechanism contributes to improved model performance and provides insight into the effectiveness of this architectural component.
@@ H.4 The role of positional encodings
Overall, the inclusion of any form of positional encoding significantly improves generation validity compared to the baseline with no PE. Among the methods tested, directed encodings like RRWP and magnetic Laplacian-based variants demonstrate particularly strong performance. These encodings consistently enhance graph validity and reduce MMD ratio in both datasets. Notably, the MagLap variant with multiple potentials ($Q=10$) achieves the best results in terms of validity and structural fidelity, although resulting in a higher computational cost.
Interestingly, although classical Laplacian encodings (Lap) also play a role in increasing the V.U.N. ratio, they obtain worse results than the directed counterparts, emphasizing that for directed generation tasks, the choice of positional encoding is not merely a design detail but a relevant architectural decision that affects model success.
@@
H.5 ER vs DAG performance
@@ H.6 Scalability experiments
SBM RRWP - Double 0.0083 $\pm$ 0.0024 0.0081 $\pm$ 0.0025 0.1789 $\pm$ 0.0107 0.0084 $\pm$ 0.0014 0.0084 $\pm$ 0.0067 27.1 $\pm$ 2.3 0.0 $\pm$ 0.0 100 $\pm$ 0.0 100 $\pm$ 0.0 0.0 $\pm$ 0.0
RRWP - Dual 0.0038 $\pm$ 0.0011 0.0037 $\pm$ 0.0012 0.0492 $\pm$ 0.0041 0.0035 $\pm$ 0.0007 0.0015 $\pm$ 0.0004 1.4 $\pm$ 0.2 87.0 $\pm$ 5.1 100 $\pm$ 0.0 100 $\pm$ 0.0 87.0 $\pm$ 5.1
MagLap - Double 0.0081 $\pm$ 0.0025 0.0079 $\pm$ 0.0020 0.1799 $\pm$ 0.0097 0.0307 $\pm$ 0.0052 0.0723 $\pm$ 0.0156 24.8 $\pm$ 4.6 6.0 $\pm$ 4.9 100 $\pm$ 0.0 100 $\pm$ 0.0 6.0 $\pm$ 4.9
@@ H.7 Conditional generation
H.8 Impact of sampling optimization using discrete flow matching
@@ I.1 Graph iterative refinement methods
$\mathbf{p}^{\theta}_{1|t}(\cdot|G_{t})=\left(\left(p_{1\mid t}^{\theta,(n)}(x_{1}^{(n)}\mid G_{t})\right)_{1\leq n\leq N},\left(p_{1\mid t}^{\theta,(i,j)}(e_{1}^{(i,j)}\mid G_{t})\right)_{1\leq i\neq j\leq N}\right),$ (38)
which predicts categorical distributions over nodes and edges given a noised graph $G_{t}$. In practice, $\mathbf{p}^{\theta}_{1|t}(\cdot|G_{t})$ is parameterized using a graph transformer such as the one from Dwivedi & Bresson (2021), which allows for expressive modeling of complex graph structures. The network is trained using a cross-entropy loss applied independently to each node and each edge:
@@ I.3 Discrete diffusion for graph generation
Among discrete diffusion-based models for undirected graphs, DiGress (Vignac et al., 2023a), grounded in the structured discrete diffusion framework (Austin et al., 2021), has been particularly influential. This approach mostly differs from the DFM-based formulation described in Section I.2, as it operates in discrete time, with the denoising process modeled as a Discrete-Time Markov Chain (DTMC), in contrast to the continuous-time formulation used in DFM.
@@
where $\alpha^{t}$ transitions from 1 to 0 with $t$ according to the popular cosine scheduling (Nichol & Dhariwal, 2021). The vectors $\mathbf{1}_{X}\in\{1\}^{X}$ and $\mathbf{1}_{E}\in\{1\}^{E+1}$ are filled with ones, and $\mathbf{m}_{X}\in\Delta^{X}$ and $\mathbf{m}_{E}\in\Delta^{E+1}$ are vectors filled with the marginal node and edge distributions, respectively [^2] We denote the probability simplex of the state-space of cardinality $Z$ by $\Delta^{Z}$.
@@
Although Directo shows strong performance in the generation of directed graphs, several avenues for future work could further enhance its scalability, controllability, and applicability.

### v3

@@ Appendix B Positional Encodings
Directed Laplacian (DirLap)
The directed Laplacian Chung (2005) is defined as
$\bm{L}=\bm{I}-\frac{1}{2}\left(\bm{\Phi}^{\frac{1}{2}}\bm{P}\bm{\Phi}^{-\frac{1}{2}}+\bm{\Phi}^{-\frac{1}{2}}\bm{P}\bm{\Phi}^{\frac{1}{2}}\right),$ (31)
where $\bm{I}$ is the identity matrix, $\bm{P}$ is the transition matrix of the directed graph, and $\bm{\Phi}$ is a diagonal matrix containing the Perron vector of $\bm{P}$. Depending on the chosen walk type, $\bm{P}$ may correspond to a standard random walk, a lazy random walk, or a random walk with teleportation (PageRank). This construction results in real eigenvalues and eigenvectors and, as a consequence, it cannot jointly encode magnitude and phase information, limiting its ability to capture directional patterns.
@@ Appendix C Sampling Optimization in Directo
Together, these sampling parameters constitute a flexible, principled toolkit for tailoring the sampling procedure to diverse graph domains and structural requirements. In our case, we optimize these parameters at sampling time to improve the performance of Directo. In particular, we individually optimize each of them for the main results, and perform the final sampling using the best performing parameters in terms of V.U.N (a detailed description is available in Appendix G.2). and ratio performance. A study of the effect of these parameters on sampling performance can be found in H.9.
@@ F.1 Validity metrics
Given a graph $G=(V,E)$ with $n=|V|$ nodes and $m=|E|$ edges, we want to determine whether the graph likely originates from an Erdős–Rényi model $G(n,p)$ with edge probability $p$ using a Wald test (Held and Sabanés Bové, 2013). For that, we follow the following steps distinguishing the cases in which the graphs are DAGs or not:
@@ G.2 Training setup
To perform sampling optimization, we conducted a targeted search over the three key hyperparameters: the time distortion function, the target guidance factor ($\omega$), and the stochasticity coefficient ($\eta$). Each hyperparameter was varied independently, while the others were held at their default values (identity distortion and $\omega=\eta=0$). For computational efficiency, searches were performed using 100 sampling steps for each configuration with five independent sampling runs to ensure robustness. An exception was made for the TPU Tiles dataset, where only a single sampling run was performed per configuration due to the large size of the dataset and some individual graphs.
@@
ER-DAG SBM TPU Tiles Visual Genome
Encoding Distortion $\bm{\omega}$ $\bm{\eta}$ Distortion $\bm{\omega}$ $\bm{\eta}$ Distortion $\bm{\omega}$ $\bm{\eta}$ Distortion $\bm{\omega}$ $\bm{\eta}$
Baseline Polydec 0.3 10 Revcos 0 0 Polydec 0 0 Cos 0.1 0
RRWP Polydec 0.1 25 Polyinc 0.1 50 Polyinc 0.3 10 Polydec 0.3 10
MagLap Polydec 0 100 Polydec 0.01 50 Polydec 0.5 100 Polydec 0.1 50
@@ G.3 Resources and runtime
Table 8: Training and sampling runtimes, and VRAM usage, across datasets and models.
Model # Train Graphs Train Time (h) # Sampled Sample Time (min/sample) VRAM (GB)
@@
MLE 128 0.001 40 0.02 0.15
D-VAE 128 33.8 40 1.2 9.07
LayerDAG 128 0.2 40 0.1 1.21
DiGress 128 22 40 0.2 5.21
DeFoG 128 13.6 40 0.1 5.71
Directo-DD (RRWP) 128 14.5 40 0.2 5.21
Directo-DD (MagLap Q=1Q{=}1) 128 18.2 40 0.4 5.21
Directo-DD (MagLap Q=5Q{=}5) 128 20 40 1.2 5.24
Directo-DD (MagLap Q=10Q{=}10) 128 26.5 40 2.5 5.30
Directo (RRWP) 128 13.5 40 0.3 5.72
Directo (MagLap Q=1Q{=}1) 128 14.4 40 0.5 5.72
Directo (MagLap Q=5Q{=}5) 128 16 40 0.8 5.75
Directo (MagLap Q=10Q{=}10) 128 25 40 1.5 5.71
@@
MLE 128 0.002 40 0.01 0.15
DiGress 128 15 40 0.2 26.81
DeFoG 128 23.7 40 0.3 28.46
Directo-DD (RRWP) 128 21.3 40 0.4 26.83
Directo-DD (MagLap Q=1Q{=}1) 128 20.4 40 1.3 26.81
Directo-DD (MagLap Q=5Q{=}5) 128 21 40 1.8 26.85
Directo-DD (MagLap Q=10Q{=}10) 128 20 40 3.1 26.81
Directo (RRWP) 128 13.7 40 0.5 28.34
Directo (MagLap Q=1Q{=}1) 128 16 40 2.1 28.45
Directo (MagLap Q=5Q{=}5) 128 19 40 5.1 28.49
Directo (MagLap Q=10Q{=}10) 128 20 40 6.0 28.46
@@
MLE 5040 0.26 40 0.01 0.15
D-VAE 5040 OOM OOM OOM OOM
LayerDAG 5040 1 40 0.5 1.75
DiGress 5040 26.3 40 1.9 30.56
DeFoG 5040 20.3 40 1.1 30.58
Directo-DD (RRWP) 5040 24.3 40 2.2 30.62
Directo-DD (MagLap Q=1Q{=}1) 5040 25 40 2.4 30.56
Directo-DD (MagLap Q=5Q{=}5) 5040 31 40 2.9 30.60
Directo (RRWP) 5040 21 40 1.1 30.60
Directo (MagLap Q=1Q{=}1) 5040 30 40 2.8 30.58
Directo (MagLap Q=5Q{=}5) 5040 30 40 3.2 30.61
@@
MLE 203 0.02 63 0.01 0.15
DiGress 203 11 63 1.0 11.03
DeFoG 203 13.6 63 0.9 11.22
Directo-DD (RRWP) 203 8 63 1.0 11.25
Directo-DD (MagLap Q=1Q{=}1) 203 9 63 1.4 11.03
Directo-DD (MagLap Q=5Q{=}5) 203 12 63 1.5 11.06
Directo (RRWP) 203 9 63 1.1 11.23
Directo (MagLap Q=1Q{=}1) 203 8 63 1.2 11.22
Directo (MagLap Q=5Q{=}5) 203 10 63 1.3 11.25
@@ H.2 Extended results for real-world datasets
Table 11: Directed graph generation performance on node-label distributional alignment metrics for the two real-world datasets. Results are presented as mean $\pm$ standard deviation across five different sampling runs.
@@ H.3 The role of dual attention
Table 16 presents detailed ablation results highlighting the impact of the dual attention mechanism in our method. The results are presented for two datasets: ER-DAG and SBM. In particular, we observe that, across both datasets, the inclusion of the dual attention mechanism consistently improves the validity of the generated graphs and reduces structural discrepancies as measured by clustering, spectral, and wavelet distances. In the ER-DAG setting, models with dual attention (e.g., RRWP-Dual and MagLap-Dual) significantly outperform their non-dual counterparts in validity, achieving up to 94% validity with RRWP-Dual and 91% with MagLap-Dual, compared to just 67.5% and 74%, respectively. These models also demonstrate better structural fidelity, particularly in clustering and spectral metrics. A similar trend is observed in the SBM dataset, where dual attention boosts validity (e.g., RRWP-Dual: 87%, MagLap-Dual: 77%). These results demonstrate how incorporating bidirectional information flow into the attention mechanism contributes to improved model performance and provides insight into the effectiveness of this architectural component.
@@
RRWP 0.0026 $\pm$ 0.0019 0.0022 $\pm$ 0.0016 0.0813 $\pm$ 0.0058 0.0048 $\pm$ 0.0002 0.0110 $\pm$ 0.0019 4.3 $\pm$ 0.8 37 $\pm$ 6.6 100 $\pm$ 0.0 100 $\pm$ 0.0 37.0 $\pm$ 6.6
RRWP - Dual 0.0031 $\pm$ 0.0014 0.0028 $\pm$ 0.0013 0.0594 $\pm$ 0.0027 0.0035 $\pm$ 0.0004 0.0027 $\pm$ 0.0009 1.8 $\pm$ 0.5 99.5 $\pm$ 1.0 100 $\pm$ 0.0 100 $\pm$ 0.0 99.5 $\pm$ 1.0
MagLap 0.0032 $\pm$ 0.0014 0.0028 $\pm$ 0.0013 0.1229 $\pm$ 0.0097 0.0068 $\pm$ 0.0010 0.0198 $\pm$ 0.0016 7.3 $\pm$ 0.7 31.5 $\pm$ 2.0 100 $\pm$ 0.0 100 $\pm$ 0.0 31.5 $\pm$ 2.0
@@ H.4 The role of positional encodings
Overall, the inclusion of any form of positional encoding significantly improves generation validity compared to the baseline with no PE, as expected. Among the methods tested, directed encodings like RRWP and magnetic Laplacian-based variants demonstrate particularly strong performance. These encodings consistently enhance graph validity and reduce MMD ratio in both datasets. Notably, the MagLap variant with multiple potentials ($Q=10$) achieves the best results in terms of validity and structural fidelity, although resulting in a higher computational cost.
@@
Interestingly, although classical Laplacian encodings (Lap and DirLap) also play a role in increasing the V.U.N. and overall ratio, they obtain worse results than the variants that explicitly encode position and directionality. This emphasizes that, for directed generation tasks, the choice of positional encoding is not merely a design detail, but a relevant architectural decision that affects model success.
@@
H.5 Ablation of other model components
To evaluate the importance of other parts of our proposed architecture, we have performed an ablation on the role of FiLM layers to enrich the attention with edge information, the use of the softmax in the aggregation of the source and target attention maps, and the gated residual connection for the node update. The results in Table support the addition of these model components, since it results in improved model performance.
Table 16: Ablation results for different model components. All trainings use dual attention and RRWP as the positional encoding.
Dataset Model Ratio $\downarrow$ Valid $\uparrow$ Unique $\uparrow$ Novelty $\uparrow$ V.U.N. $\uparrow$
ER-DAG Original 1.7 $\pm$ 0.1 94 $\pm$ 1.0 100 $\pm$ 0.0 100 $\pm$ 0.0 94 $\pm$ 1.0
No FILM 6.9 $\pm$ 0.6 4 $\pm$ 3.0 100 $\pm$ 0.0 100 $\pm$ 0.0 4 $\pm$ 3.0
No softmax 1.7 $\pm$ 0.2 74.5 $\pm$ 1.9 100 $\pm$ 0.0 100 $\pm$ 0.0 74.5 $\pm$ 1.9
No FILM + No softmax 6.3 $\pm$ 0.2 0.0 $\pm$ 0.0 100 $\pm$ 0.0 100 $\pm$ 0.0 0.0 $\pm$ 0.0
No gated residual connection 2.5 $\pm$ 0.2 66.5 $\pm$ 6.0 100 $\pm$ 0.0 100 $\pm$ 0.0 66.5 $\pm$ 6.0
SBM Original 1.8 $\pm$ 0.5 99.5 $\pm$ 3.7 100 $\pm$ 0.0 100 $\pm$ 0.0 99.5 $\pm$ 3.7
No FILM 15.8 $\pm$ 3.3 0.0 $\pm$ 0.0 100 $\pm$ 0.0 100 $\pm$ 0.0 0.0 $\pm$ 0.0
No softmax 2.4 $\pm$ 1.4 89.0 $\pm$ 3.7 100 $\pm$ 0.0 100 $\pm$ 0.0 89.0 $\pm$ 3.7
No FILM + No softmax 16.7 $\pm$ 2.8 0.0 $\pm$ 0.0 100 $\pm$ 0.0 100 $\pm$ 0.0 0.0 $\pm$ 0.0
No gated residual connection 2.7 $\pm$ 1.4 91.5 $\pm$ 1.7 100 $\pm$ 0.0 100 $\pm$ 0.0 91.5 $\pm$ 1.7
H.6 ER vs DAG performance
@@ H.7 Scalability experiments
SBM RRWP - Double 0.0085 $\pm$ 0.0025 0.0081 $\pm$ 0.0026 0.1755 $\pm$ 0.0135 0.0078 $\pm$ 0.0009 0.0816 $\pm$ 0.0075 26.3 $\pm$ 2.5 0.0 $\pm$ 0.0 100 $\pm$ 0.0 100 $\pm$ 0.0 0.0 $\pm$ 0.0
RRWP - Dual 0.0031 $\pm$ 0.0014 0.0028 $\pm$ 0.0013 0.0594 $\pm$ 0.0027 0.0035 $\pm$ 0.0004 0.0027 $\pm$ 0.0009 1.8 $\pm$ 0.5 99.5 $\pm$ 1.0 100 $\pm$ 0.0 100 $\pm$ 0.0 99.5 $\pm$ 1.0
MagLap - Double 0.0086 $\pm$ 0.0023 0.0083 $\pm$ 0.0024 0.1878 $\pm$ 0.0221 0.0152 $\pm$ 0.0044 0.0376 $\pm$ 0.0037 14.2 $\pm$ 1.5 8.0 $\pm$ 7.5 100 $\pm$ 0.0 100 $\pm$ 0.0 8.0 $\pm$ 7.5
@@ H.8 Conditional generation
H.9 Impact of sampling optimization using discrete flow matching
@@ I.1 Graph iterative refinement methods
$\bm{p}^{\theta}_{1|t}(\cdot|G_{t})=\left(\left(p_{1\mid t}^{\theta,(n)}(x_{1}^{(n)}\mid G_{t})\right)_{1\leq n\leq N},\left(p_{1\mid t}^{\theta,(i,j)}(e_{1}^{(i,j)}\mid G_{t})\right)_{1\leq i\neq j\leq N}\right),$ (39)
which predicts categorical distributions over nodes and edges given a noised graph $G_{t}$. In practice, $\mathbf{p}^{\theta}_{1|t}(\cdot|G_{t})$ is parameterized using a graph transformer such as the one from Dwivedi and Bresson (2021), which allows for expressive modeling of complex graph structures. The network is trained using a cross-entropy loss applied independently to each node and each edge:
@@ I.3 Discrete diffusion for graph generation
Among discrete diffusion-based models for undirected graphs, DiGress (Vignac et al., 2023a), grounded in the structured discrete diffusion framework (Austin et al., 2021a), has been particularly influential. This approach mostly differs from the DFM-based formulation described in Section I.2, as it operates in discrete time, with the denoising process modeled as a Discrete-Time Markov Chain (DTMC), in contrast to the continuous-time formulation used in DFM.
@@
where $\alpha^{t}$ transitions from 1 to 0 with $t$ according to the popular cosine scheduling (Nichol and Dhariwal, 2021). The vectors $\mathbf{1}_{X}\in\{1\}^{X}$ and $\mathbf{1}_{E}\in\{1\}^{E+1}$ are filled with ones, and $\mathbf{m}_{X}\in\Delta^{X}$ and $\mathbf{m}_{E}\in\Delta^{E+1}$ are vectors filled with the marginal node and edge distributions, respectively [^2] We denote the probability simplex of the state-space of cardinality $Z$ by $\Delta^{Z}$.
@@
Although Directo shows strong performance in the generation of directed graphs, it suffers from inheren lmitiations, that open several avenues for future work. These are mostly addressed at further enhancing its scalability, controllability, and applicability.